OpenStudy (anonymous):

2 +3) + 4 = blank + (2+4) please

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yay thank you

OpenStudy (karatechopper):

do u wanna know how to solve it?

OpenStudy (anonymous):

yeah

OpenStudy (karatechopper):

ok (2+3)+4=n+ (2+4) i put the n in there as a blank u will under stand in like 5th or 6th grade the variable stuff. 5+4=n+6 9=n+6 9-6=n+6-6 3=n